# Euclidean Distance Matrix Completion via Asymmetric Projected Gradient Descent

### v1

@@ III-D Iteration Trajectory
Figure 5: The phase transition measured by EDM recovery rate of APGD when varying protein type and Bernoulli sample rate $p$, while fixing $r=3$. Detailed parameters for each protein are listed in Table I. The results are averaged over $5$ independent trials.
A special instance of experiments in Section III-B is investigated in greater detail. We consider $(n,p)=(1500,0.1)$, $(n,p)=(1500,0.05)$, and plot the convergence trajectories measured by (16) in Fig. 4. When $p=0.1$, 96% of the trials successfully recover the EDM. The corresponding trajectories of $r^{k}$, $g^{k}_{1}$, and $g^{k}_{2}$ are presented in Fig. 4(a), (b), and (c) respectively, where exact linear convergence is observed after a few iterations, as indicated by the gray lines. Moreover, the pseudo gradient and $g^{k}_{2}$ demonstrated nearly identical behavior. When $p=0.05$, 47% of the recordings are categorized to be successful [^12] This clearly highlights the performance degradation caused by imposing the lower safeguard $\eta_{\min}$. Also, the succeeded instances of $g^{k}_{1}$ and $g^{k}_{2}$ when $p=0.05$ resemble those in Fig. 4(b), (c), and thus omitted here., with their $r^{k}$ shown in Fig. 4(d). Intriguing phenomena emerge when analyzing the other failed $53\%$, as illustrated in Fig. 4(e), (f), where $g^{k}_{1}$ and $g^{k}_{2}$ no longer coincide. In about $17\%$ of all trials, the pseudo gradient tends to converge, whereas the corresponding $g^{k}_{2}$ does not. This observation might uncover a potential factor underlying the degraded numerical performance of APGD, i.e., the consistency between pseudo gradient and the real gradient of (4a) (up to trivial rescale) demands an unreasonably large number of samples to be reliably ensured.
@@ III-E Protein Test
Phase transition tests on nine proteins downloaded from the Protein Data Bank[68] are presented in Fig. 5 to evaluate the real-world performance of APGD, as well as the impact of condition number $\kappa$ and coherence parameter $\mu$. Similar experiments but under the unit ball sample model and challenging noisy observations settings can be found in[9, Sec. 5.3]. Prior to processing, solvent molecules and chelated metal ions were removed from the datasets. The resulting parameters $n$, $\mu$, and $\kappa$ are reported in Table I.
Non-monotonic trends in sample complexity $p$ w.r.t. $n$ are observed, e.g., although the atom number of 1I7W is approximately three times larger than that of 1KDH, the phase transition for 1KDH occurs at a significantly lower sampling ratio. This phenomenon is attributed to the large condition number associated with the 1I7W configuration, which both suppresses the BB stepsize and increases the sample complexity required for global recovery, as predicted by Theorem II.1. Furthermore, the performance of Algorithm 2 appears sensitive to the product $\mu\kappa$, since the curves of 1AX8 versus 1RGS, and 1TIM versus 1TOA exhibit similar “reversal” [^13] That is, the phase transition for a protein with larger $n$ occurs later than that for a protein with smaller $n$., due to small changes in their corresponding $\mu\kappa$ values.
@@ IV Conclusion and Discussion
  • Since the EDM basis shares a similar support pattern with $\mathbf{e}_{i}\mathbf{e}_{j}^{T}$, [20] introduced the incoherence w.r.t. $\boldsymbol{\omega}_{\boldsymbol{\alpha}}$ and $\boldsymbol{\nu}_{\boldsymbol{\alpha}}$ to measure the concentration of information in $\mathbf{G}^{\star}$. While such definition, c.f.,[41, Def. 2], can be deduced from classical incoherence assumption up to trivial rescale by resorting to[23, Lemma A.4], this does not fully account for the observed performance gap between trace minimization and VGD applied to (1). In the context of Hankel Matrix Completion (HMC), incoherence is satisfied if the minimum wraparound distance between spikes remains non-vanishing[72, Sec. III-A], and both EMaC (convex approach to HMC) and PGD-like algorithms[45][73] are observed to succeed even without strict separation. But for EDMC, which property of the point set govern the problem’s tractability?
@@ ACKNOWLEDGMENT
The authors would like to thank Abiy Tasissa for insightful discussion and help in proofreading the manuscript.
@@
$\|\Delta_{1}\|_{F}^{2}:$ $=\|\mathcal{P}_{I}(\hat{\mathbf{P}}-\mathbf{P}^{\star}\boldmath{\psi}^{\star})\|_{F}^{2}\overset{(i)}{\leq}\|\hat{\mathbf{P}}-\mathbf{P}^{\star}\boldmath{\psi}^{\star}\|_{F}^{2}$
@@
$\overset{(iv)}{\leq}\frac{r\epsilon^{2}\sigma_{1}^{\star 2}}{2(\sqrt{2}-1)\sigma_{r}^{\star}}\overset{(v)}{\lesssim}\frac{\sigma_{r}^{\star}}{c_{c}},$
where $(i)$ uses the convexity of the set $\{\mathbf{P}:\|\mathbf{P}\|_{2,\infty}\leq\sqrt{\frac{c_{I}\mu r\sigma_{1}^{\star}}{n}}\}$, see also [35, Lemma 11]. $(ii)$ comes from [54, Lemma 6], and $(iii)$ uses $\|\mathbf{A}\|_{F}^{2}\leq r\|\mathbf{A}\|^{2}$ if $\mathrm{rank}(\mathbf{A})=r$. $(iv)$ comes from Lemma C.6, and holds with probability at least $1-n^{1-\beta}$ given $p\gtrsim\frac{C_{\beta}\mu^{2}r^{2}\log n}{\epsilon^{2}n}$. Finally $(v)$ by setting $\epsilon^{2}\leq\frac{1}{C_{\epsilon}\kappa^{2}r}$, where $C_{\epsilon}\gtrsim c_{c}$, concluding the proof.∎
@@
where $\mathcal{P}_{\mathbf{U}}=\mathbf{U}^{\star}\mathbf{U}^{\star T}$, and $\mathbf{W}_{1}\in\mathbb{R}^{n\times r}$ are arbitrary. Let ^$\hat{\nabla}f(\mathbf{P}):=\frac{2}{p}\mathcal{R}_{\Omega}(\mathbf{P}\mathbf{P}^{T}-\mathbf{G}^{\star})\mathbf{P}$ denote the pseudo gradient. We will also frequently use Lemma C.7 without explicitly referring to it.
B-A Local Contraction With Re-sampling
It is meanwhile straightforward to prove the local contraction property by following the roadmap in, e.g., [35, App. D][45, App. C][73], given the following claim holds for the moment.
@@ B-A Local Contraction With Re-sampling
^$\langle\hat{\nabla}f(\mathbf{P}),\Delta\rangle\geq\alpha\|\Delta\|_{F}^{2},$ (19a)
^$\|\hat{\nabla}f(\mathbf{P})\|_{F}^{2}\leq\rho\|\Delta\|_{F}^{2},$ (19b)
^$\langle\hat{\nabla}f(\mathbf{P}),\Delta\rangle\geq\frac{\alpha}{2}\|\Delta\|_{F}^{2}+\frac{\alpha}{2\rho}\|\hat{\nabla}f(\mathbf{P})\|_{F}^{2},$ (19c)
@@ B-B Restricted Strong Convexity: (19a)
$\overset{(i)}{\geq}\frac{5}{8}\overbrace{\|\Delta\bar{\mathbf{P}}^{\star T}+\bar{\mathbf{P}}^{\star}\Delta^{T}\|_{F}^{2}}^{a^{2}}-4\overbrace{\|\Delta\Delta^{T}\|_{F}^{2}}^{b^{2}},$
@@
$A_{2}\leq\frac{2}{5}\sigma_{r}^{\star}\|\Delta\|_{F}^{2}$ (28)
@@ B-C Restricted Smoothness: (19b)
Notice that ^$\|\hat{\nabla}f(\mathbf{P})\|_{F}^{2}=|\sup_{\|\mathbf{Z}\|_{F}^{2}=1}\langle\hat{\nabla}f(\mathbf{P}),\mathbf{Z}\rangle|^{2}$, and $A_{3}:=|\langle\hat{\nabla}f(\mathbf{P}),\mathbf{Z}\rangle|^{2}$ can be separated as in (23). Define $\mathbf{Z}_{\Delta}:=\mathbf{Z}\Delta^{T}+\Delta\mathbf{Z}^{T}$, and $\mathbf{Z}_{P^{\star}}=\mathbf{Z}\bar{\mathbf{P}}^{\star T}+\bar{\mathbf{P}}^{\star}\mathbf{Z}^{T}\in\mathbb{T}$, we have
@@
and $\sqrt{\frac{1}{p}\|\mathcal{P}_{\Omega}\mathcal{P}_{\mathbb{T}}(\mathbf{Z}_{P^{\star}})\|_{F}^{2}}\overset{(c)}{\leq}\sqrt{\frac{11}{10}\|\mathbf{Z}_{P^{\star}}\|_{F}^{2}}\leq\frac{\sqrt{110\sigma_{1}^{\star}}}{5}\|\mathbf{Z}\|_{F}$. Where $(a)$ from Lemma C.5 and then use the RIC assumption on $\mathcal{B}$, $(b)$ by forcing $p\gtrsim\frac{C_{\beta}\mu^{2}r^{2}\log n}{n}$, and $(c)$ from Lemma C.4 with $\epsilon=0.1$ and $p\gtrsim\frac{C_{\beta}\mu r\log n}{\epsilon^{2}n}$. We infer current bound on $\frac{1}{p}\|\mathcal{Q}_{\Omega}(\Delta\Delta^{T})\|_{F}^{2}$, i.e., [22, Lemma B.2] or Lemma C.5, is not sharp up to $\sqrt{n}$, but we cannot amend it. If this estimate can be shown to be as good as the random graph lemma[33, Lemma 7.1], then one will immediately obtain the local convergence mechanism, i.e., the contraction speed measured by $\|\Delta\|_{F}$, when optimizing the s-stress function.
@@
^$\|\hat{\nabla}f(\mathbf{P})\|_{F}^{2}\leq 84\sigma_{1}^{\star 2}c_{I}\mu r\|\Delta\|_{F}^{2},$ (35)
@@ Appendix C Supporting Lemmas
$\mathbb{E}\mathbf{S}_{k}=\mathbf{0},\,\|\mathbf{S}_{k}\|\leq B\,\,\text{almostsurely.}$.
@@
The random graph lemma, see also[33, Lemma 7.1]. If $p\geq\frac{C_{g}\beta\log n}{n}$ for some $C_{g}>0$, $\beta>1$,
@@
The local Restricted Isometry Property (RIP) of $\mathcal{R}_{\Omega}$. If $\|\mathbf{U}^{\star}\|_{2,\infty}^{2}\leq\frac{\mu r}{n}$, then for $\epsilon\geq\sqrt{\frac{C_{r}\beta(\mu r)^{2}\log n}{np}}$
@@
$\frac{1}{p}\|\mathcal{Q}_{\Omega}$ $(\Delta\Delta^{T})\|_{F}^{2}$
$\leq\left[(8n+\sqrt{\frac{c\beta n\log n}{p}})\|\Delta\|_{2,\infty}^{2}+8\|\Delta\|_{F}^{2}\right]\|\Delta\|_{F}^{2},$
@@ Appendix D Proof of Auxiliary Lemmas
The proof of Lemma B.1 and B.2 is partly inspired by [70, Lemma 8][71, Lemma 22], and [61, Thm. 4.1], respectively.
@@
=$\overset{(i)}{=}\sup_{\mathbf{x},\mathbf{y}\in\mathbb{S}^{n-1}}\bigg{\langle}\frac{1}{p}\mathcal{P}_{\Omega}(\mathbf{H}_{\mathbf{1}})-\mathbf{H}_{\mathbf{1}},g(\mathbf{AB}^{T})\circ\mathbf{xy}^{T}\bigg{\rangle}$
@@
Same bound holds for $I_{22}$. By [69, Lemma 5], $I_{23}$ can be decomposed as
$I_{23}=\bigg{\|}(\sum_{k=1}^{r}\mathbf{A}_{.,k}\mathbf{B}_{.,k}^{T})\circ\mathbf{xy}^{T}\bigg{\|}_{*}=\bigg{\|}\sum_{k=1}^{r}(\mathbf{A}_{.,k}\circ\mathbf{x})(\mathbf{B}_{.,k}\circ\mathbf{y})^{T}\bigg{\|}_{*}$
$\leq\sum_{k=1}^{r}\|(\mathbf{A}_{.,k}\circ\mathbf{x})(\mathbf{B}_{.,k}\circ\mathbf{y})^{T}\|_{*}\leq\sum_{k=1}^{r}\|\mathbf{A}_{.,k}\circ\mathbf{x}\|_{2}\|\mathbf{B}_{.,k}\circ\mathbf{y}\|_{2}$
$\overset{(i)}{\leq}\sqrt{\sum_{i=1}^{n}\mathbf{x}_{i}^{2}\sum_{k_{1}=1}^{r}\mathbf{A}_{ik_{1}}^{2}}\sqrt{\sum_{j=1}^{n}\mathbf{y}_{j}^{2}\sum_{k_{2}=1}^{r}\mathbf{B}_{jk_{2}}^{2}}$
$\leq\|\mathbf{A}\|_{2,\infty}\|\mathbf{B}\|_{2,\infty},$ (37)
where $(i)$ from Cauchy-Schwarz. Since $\|\mathbf{AB}^{T}\|_{\infty}\leq\|\mathbf{A}\|_{2,\infty}\|\mathbf{B}\|_{2,\infty}$, we have
@@
Unfortunately, same strategy as in App. D-A does not applies to Lemma B.2, due to the asymmetry structure of $\mathcal{R}_{\Omega}$. Uniform bound presents here largely dues to the fixed, incoherent matrix $\mathbf{B}$. Let $\mathbf{X}=\mathbf{A}\boldsymbol{R}^{T}\mathbf{B}^{T}$, we need to control $T_{0}=\|(\frac{1}{p}\mathcal{R}_{\Omega}^{*}-\mathcal{I})\mathbf{X}\|$, where $\mathbf{X}=\mathbf{JXJ}$ always holds. Notice that by using the structure of $g^{*}(\mathbf{D}):=2(\mathrm{diag}(\mathbf{D}\mathbf{1})-\mathbf{D})$, $T_{0}$ can be reduced to
@@
$\overset{(ii)}{\leq}\underbrace{\bigg{\|}(\frac{1}{p}\mathcal{P}_{\Omega}\mathbf{X}-\mathrm{Od}(\mathbf{X}))\mathbf{1}\bigg{\|}_{\infty}}_{T_{1}}+\underbrace{\bigg{\|}\frac{1}{p}\mathcal{P}_{\Omega}\mathbf{X}-\mathrm{Od}(\mathbf{X})\bigg{\|}}_{T_{2}}$
@@
Bound on $T_{2}$: This follows from similar argument as in[71, Lemma 22], thus holds uniformly over $\mathbf{A}$, $\mathbf{B}$, and $\boldsymbol{R}$. By using the same trick as in App. D-A, one can show
$T_{2}$ $=\sup_{\mathbf{x},\mathbf{y}\in\mathcal{S}^{n-1}}\bigg{\langle}\frac{1}{p}\mathcal{P}_{\Omega}(\mathbf{H}_{\mathbf{1}})-\mathbf{H}_{\mathbf{1}},\mathrm{Od}(\mathbf{A}\bar{\mathbf{B}}^{T})\circ\mathbf{xy}^{T}\bigg{\rangle}$
@@
and the same upper bound in (37) applies to $T_{31}$. For $T_{32}$, first recall the column blocking structure, we have
@@
$\overset{(ii)}{\leq}\sqrt{\sum_{i=1}^{n}\sum_{k_{1}=1}^{r}\mathbf{A}_{ik_{1}}^{2}\mathbf{x}_{i}^{2}}\sqrt{\sum_{j=1}^{n}\sum_{k_{2}=1}^{r}\bar{\mathbf{B}}_{jk_{2}}^{2}\mathbf{y}_{j}^{2}}$
@@
where $(i)$, $(ii)$ use Cauchy-Schwarz, $(iii)$ from same argument as in (37). Thus $T_{3}\leq 2\|\mathbf{A}\|_{2,\infty}\|\mathbf{B}\|_{2,\infty}$. By substituting (39) into (38), we conclude the first part of the proof.
@@
We first separate $\mathbf{A}$ in $T_{11}$, inspired by the following trick[61, Thm. 4.1]
@@
$\bar{T}_{11}$ $\leq\frac{C\beta\log n\|\mathbf{B}\|_{2,\infty}}{p}+\sqrt{\frac{C\beta n\log n\|\mathbf{B}\|_{2,\infty}^{2}}{p}}$
$\leq\sqrt{\frac{C\beta n\log n}{p}}\|\mathbf{B}\|_{2,\infty},$ (41)
@@
The original proof[75, App. B-D] has some flaws, but we have amended it here. Notice that by (12) and Cauchy-Schwarz, we have
@@
$\leq(8\left\|\sum_{\boldsymbol{\alpha}\in\mathbb{I}}\mathbf{B}_{\boldsymbol{\alpha}}\right\|+8n)\|\mathbf{x}\|_{2}^{2}+8\|\mathbf{x}\|_{1}^{2}$ (42)

### v2

@@ III-D Iteration Trajectory
A special instance of experiments in Section III-B is investigated in greater detail. We consider $(n,p)=(1500,0.1)$, $(n,p)=(1500,0.05)$, and plot the convergence trajectories measured by (19) in Fig. 4. When $p=0.1$, 96% of the trials successfully recover the EDM. The corresponding trajectories of $r^{k}$, $g^{k}_{1}$, and $g^{k}_{2}$ are presented in Fig. 4(a), (b), and (c) respectively, where exact linear convergence is observed after a few iterations, as indicated by the gray lines. Moreover, the pseudo gradient and $g^{k}_{2}$ demonstrated nearly identical behavior. When $p=0.05$, 47% of the recordings are categorized to be successful, with their $r^{k}$ shown in Fig. 4(d). Compared with Fig. 2(h), this clearly highlights the performance degradation caused by imposing the lower safeguard $\eta_{\min}$. Also, the succeeded instances of $g^{k}_{1}$ and $g^{k}_{2}$ when $p=0.05$ resemble those in Fig. 4(b), (c), and thus omitted here. Intriguing phenomena emerge when analyzing the other failed $53\%$, as illustrated in Fig. 4(e), (f), where $g^{k}_{1}$ and $g^{k}_{2}$ no longer coincide. In about $17\%$ of all trials, the pseudo gradient tends to converge, whereas the corresponding $g^{k}_{2}$ does not. This observation might uncover a potential factor underlying the degraded numerical performance of APGD, i.e., the consistency between pseudo gradient and the real gradient of (7a) (up to trivial rescale) demands an unreasonably large number of samples to be reliably ensured.
@@ III-E Protein Test
Phase transition tests on nine proteins downloaded from the Protein Data Bank[76] are presented in Fig. 5 to evaluate the real-world performance of APGD, as well as the impact of condition number $\kappa$ and coherence parameter $\mu$. Similar experiments but under the unit ball sample model and challenging noisy observations settings can be found in[9, Sec. 5.3]. Prior to processing, solvent molecules and chelated metal ions were removed from the datasets. The resulting parameters $n$, $\mu$, and $\kappa$ are reported in Table II.
Non-monotonic trends in sample complexity $p$ w.r.t. $n$ are observed, i.e., the phase transition for a protein with larger $n$ occurs later than that for a protein with smaller $n$. For example, although the atom number of 1I7W is approximately three times larger than that of 1KDH, the phase transition for 1KDH occurs at a significantly lower sampling ratio. The large condition number associated with the 1I7W configuration suppresses the BB stepsize, and increases the sample complexity required for global recovery, as predicted by Theorem II.1. Furthermore, the performance of Algorithm 2 appears sensitive to the product $\mu\kappa$, since the curves of 1AX8 versus 1RGS, and 1TIM versus 1TOA exhibit similar “reversal”, due to small changes in their corresponding $\mu\kappa$ values.
@@ IV Conclusion and Discussion
  • Since the EDM basis shares a similar support pattern with $\mathbf{e}_{i}\mathbf{e}_{j}^{T}$, [20] introduced the incoherence w.r.t. $\boldsymbol{\omega}_{\boldsymbol{\alpha}}$ and $\boldsymbol{\nu}_{\boldsymbol{\alpha}}$ to measure the concentration of information in $\mathbf{G}^{\star}$. While such definition, c.f.,[47, Def. 2], can be deduced from classical incoherence assumption up to trivial rescale by resorting to[23, Lemma A.4], this does not fully account for the observed performance gap between trace minimization and VGD applied to (3). In the context of Hankel Matrix Completion (HMC), incoherence is satisfied if the minimum wraparound distance between spikes remains non-vanishing[40, Sec. III-A], and both EMaC (convex approach to HMC) and PGD-like algorithms[50][78] are observed to succeed even without strict separation. But for EDMC, which property of the point set govern the problem’s tractability?
@@
Figure 5: The phase transition measured by EDM recovery rate of APGD when varying protein type and Bernoulli sample rate $p$, while fixing $r=3$. Detailed parameters for each protein are listed in Table II. The results are averaged over 55 independent trials.
@@ ACKNOWLEDGMENT
The authors would like to thank Abiy Tasissa for insightful discussion and help in proofreading the manuscript. They would also like to thank the anonymous reviewers and associate editor for their constructive comments. Y. L. would like to thank Jianguo Liu for proofreading during the revision.
@@
$\|\Delta_{1}\|_{F}^{2}:$ $=\|\mathcal{P}_{I}(\hat{\mathbf{P}}-\mathbf{P}^{\star}{\psi}^{\star})\|_{F}^{2}\overset{(i)}{\leq}\|\hat{\mathbf{P}}-\mathbf{P}^{\star}{\psi}^{\star}\|_{F}^{2}$
@@
$\overset{(iv)}{\leq}\frac{r\varepsilon^{2}\sigma_{1}^{\star 2}}{2(\sqrt{2}-1)\sigma_{r}^{\star}}\overset{(v)}{\lesssim}\frac{\sigma_{r}^{\star}}{c_{c}},$
where $(i)$ uses the convexity of the set $\{\mathbf{P}:\|\mathbf{P}\|_{2,\infty}\leq\sqrt{\frac{c_{I}\mu r\sigma_{1}^{\star}}{n}}\}$, see also [38, Lemma 11]. $(ii)$ comes from [63, Lemma 6], and $(iii)$ uses $\|\mathbf{A}\|_{F}^{2}\leq r\|\mathbf{A}\|^{2}$ if $\mathrm{rank}(\mathbf{A})=r$. $(iv)$ comes from Lemma C.6, and holds with probability at least $1-n^{1-\beta}$ given $p\gtrsim\frac{C_{\beta}\mu^{2}r^{2}\log n}{\varepsilon^{2}n}$. Finally $(v)$ by setting $\varepsilon^{2}\leq\frac{1}{C\kappa^{2}r}$, where $C\gtrsim c_{c}$, concluding the proof.∎
@@
where $\mathcal{P}_{\mathbf{U}}=\mathbf{U}^{\star}\mathbf{U}^{\star T}$, and $\mathbf{W}_{1}\in\mathbb{R}^{n\times r}$ are arbitrary. Let $\hat{\nabla}f(\mathbf{P}):=\frac{2}{p}\mathcal{R}_{\Omega}(\mathbf{P}\mathbf{P}^{T}-\mathbf{G}^{\star})\mathbf{P}$ denote the pseudo gradient. We will also frequently use Lemma C.7 without explicitly referring to it.
B-A Local Contraction Without Re-sampling
It is meanwhile straightforward to prove the local contraction property by following the roadmap in, e.g., [38, App. D][50, App. C][78], given the following claim holds for the moment.
@@ B-A Local Contraction Without Re-sampling
$\langle\hat{\nabla}f(\mathbf{P}),\Delta\rangle\geq\alpha\|\Delta\|_{F}^{2},$ (22a)
$\|\hat{\nabla}f(\mathbf{P})\|_{F}^{2}\leq\rho\|\Delta\|_{F}^{2},$ (22b)
$\langle\hat{\nabla}f(\mathbf{P}),\Delta\rangle\geq\frac{\alpha}{2}\|\Delta\|_{F}^{2}+\frac{\alpha}{2\rho}\|\hat{\nabla}f(\mathbf{P})\|_{F}^{2},$ (22c)
@@ B-B Restricted Strong Convexity: (22a)
$\overset{(i)}{\geq}\frac{5}{8}\overbrace{\|\Delta\bar{\mathbf{P}}^{\star T}+\bar{\mathbf{P}}^{\star}\Delta^{T}\|_{F}^{2}}^{a^{2}}-4\overbrace{\|\Delta\Delta^{T}\|_{F}^{2}}^{b^{2}}$
@@
$A_{2}\leq\frac{2}{5}\sigma_{r}^{\star}\|\Delta\|_{F}^{2},$ (31)
@@ B-C Restricted Smoothness: (22b)
Notice that $\|\hat{\nabla}f(\mathbf{P})\|_{F}^{2}=|\sup_{\|\mathbf{Z}\|_{F}^{2}=1}\langle\hat{\nabla}f(\mathbf{P}),\mathbf{Z}\rangle|^{2}$, and $A_{3}:=|\langle\hat{\nabla}f(\mathbf{P}),\mathbf{Z}\rangle|^{2}$ can be separated as in (26). Define $\mathbf{Z}_{\Delta}:=\mathbf{Z}\Delta^{T}+\Delta\mathbf{Z}^{T}$, and $\mathbf{Z}_{P^{\star}}=\mathbf{Z}\bar{\mathbf{P}}^{\star T}+\bar{\mathbf{P}}^{\star}\mathbf{Z}^{T}\in\mathbb{T}$, we have
@@
and $\sqrt{\frac{1}{p}\|\mathcal{P}_{\Omega}\mathcal{P}_{\mathbb{T}}(\mathbf{Z}_{P^{\star}})\|_{F}^{2}}\overset{(c)}{\leq}\sqrt{\frac{11}{10}\|\mathbf{Z}_{P^{\star}}\|_{F}^{2}}\leq\frac{\sqrt{110\sigma_{1}^{\star}}}{5}\|\mathbf{Z}\|_{F}$. Where $(a)$ from Lemma C.5 and then use the RIC assumption on $\mathcal{B}$, $(b)$ by forcing $p\gtrsim\frac{C_{\beta}\mu^{2}r^{2}\log n}{n}$, and $(c)$ from Lemma C.4 with $\epsilon=0.1$ and $p\gtrsim\frac{C_{\beta}\mu r\log n}{\epsilon^{2}n}$.
@@
$\|\hat{\nabla}f(\mathbf{P})\|_{F}^{2}\leq 84\sigma_{1}^{\star 2}c_{I}\mu r\|\Delta\|_{F}^{2},$ (38)
@@ Appendix C Supporting Lemmas
$\mathbb{E}\mathbf{S}_{k}=\mathbf{0},\,\|\mathbf{S}_{k}\|\leq B\,\,\text{almost surely.}$
@@
The random graph lemma, see also[36, Lemma 7.1]. If $p\geq\frac{C_{g}\beta\log n}{n}$ for some $C_{g}>0$, $\beta>1$, $c_{g}>1$
@@
The local RIP of $\mathcal{R}_{\Omega}$. If $\|\mathbf{U}^{\star}\|_{2,\infty}^{2}\leq\frac{\mu r}{n}$, then for $\epsilon\geq\sqrt{\frac{C_{r}\beta(\mu r)^{2}\log n}{np}}$
@@
$\frac{1}{p}\|\mathcal{Q}_{\Omega}(\Delta\Delta^{T})\|_{F}^{2}\leq\left[(8n+\sqrt{\frac{c\beta n\log n}{p}})\|\Delta\|_{2,\infty}^{2}+8\|\Delta\|_{F}^{2}\right]\|\Delta\|_{F}^{2},$
@@ Appendix D Proof of Auxiliary Lemmas
The proof of Lemma B.1 and B.2 is partly inspired by [60, Lemma 8][72, Lemma 22], and [59, Thm. 4.1], respectively.
@@
$\overset{(i)}{=}\sup_{\mathbf{x},\mathbf{y}\in\mathbb{S}^{n-1}}\bigg\langle\frac{1}{p}\mathcal{P}_{\Omega}(\mathbf{H}_{\mathbf{1}})-\mathbf{H}_{\mathbf{1}},g(\mathbf{AB}^{T})\circ\mathbf{xy}^{T}\bigg\rangle$
@@
Same bound holds for $I_{22}$. By [77, Lemma 5], $I_{23}$ can be upper bounded by $\|\mathbf{A}\|_{2,\infty}\|\mathbf{B}\|_{2,\infty}$. Since $\|\mathbf{AB}^{T}\|_{\infty}\leq\|\mathbf{A}\|_{2,\infty}\|\mathbf{B}\|_{2,\infty}$, we have
@@
Unfortunately, same strategy as in App. D-A does not applies to Lemma B.2, due to the asymmetry structure of $\mathcal{R}_{\Omega}$. Uniform bound presents here dues to the fixed, incoherent matrix $\mathbf{B}$. Let $\mathbf{X}=\mathbf{A}\boldsymbol{R}^{T}\mathbf{B}^{T}$, we need to control $T_{0}=\|(\frac{1}{p}\mathcal{R}_{\Omega}^{*}-\mathcal{I})\mathbf{X}\|$, where $\mathbf{X}=\mathbf{JXJ}$ always holds. Notice that by using the structure of $g^{*}(\mathbf{D}):=2(\mathrm{diag}(\mathbf{D}\mathbf{1})-\mathbf{D})$, $T_{0}$ can be reduced to
@@
$\overset{(ii)}{\leq}\underbrace{\bigg\|(\frac{1}{p}\mathcal{P}_{\Omega}\mathbf{X}-\mathrm{Od}(\mathbf{X}))\mathbf{1}\bigg\|_{\infty}}_{T_{1}}+\underbrace{\bigg\|\frac{1}{p}\mathcal{P}_{\Omega}\mathbf{X}-\mathrm{Od}(\mathbf{X})\bigg\|}_{T_{2}},$
@@
Bound on $T_{2}$: This follows from similar argument as in[72, Lemma 22], thus holds uniformly over $\mathbf{A}$, $\mathbf{B}$, and $\boldsymbol{R}$. By using the same trick as in App. D-A, one can show
$T_{2}$ $=\sup_{\mathbf{x},\mathbf{y}\in\mathbb{S}^{n-1}}\bigg\langle\frac{1}{p}\mathcal{P}_{\Omega}(\mathbf{H}_{\mathbf{1}})-\mathbf{H}_{\mathbf{1}},\mathrm{Od}(\mathbf{A}\bar{\mathbf{B}}^{T})\circ\mathbf{xy}^{T}\bigg\rangle$
@@
and the same upper bound as $I_{23}$ applies to $T_{31}$. For $T_{32}$, first recall the column blocking structure, we have
@@
$\overset{(ii)}{\leq}\sqrt{\sum_{i=1}^{n}\mathbf{x}_{i}^{2}\sum_{k_{1}=1}^{r}\mathbf{A}_{ik_{1}}^{2}}\sqrt{\sum_{j=1}^{n}\mathbf{y}_{j}^{2}\sum_{k_{2}=1}^{r}\bar{\mathbf{B}}_{jk_{2}}^{2}}$
@@
where $(i)$, $(ii)$ use Cauchy-Schwarz, $(iii)$ from noticing $\sum_{k_{1}=1}^{r}\mathbf{A}_{ik_{1}}^{2}\leq\|\mathbf{A}\|_{2,\infty}^{2}$, $\forall i\in[n]$. Thus $T_{3}\leq 2\|\mathbf{A}\|_{2,\infty}\|\mathbf{B}\|_{2,\infty}$. By substituting (41) into (40), we conclude the first part of the proof.
@@
We first separate $\mathbf{A}$ in $T_{11}$, inspired by the following trick[59, Thm. 4.1]
@@
$\bar{T}_{11}\leq\sqrt{\frac{C\beta n\log n}{p}}\|\mathbf{B}\|_{2,\infty},$ (43)
@@
The original proof[80, App. B-D] has some flaws, but we have amended it here. Notice that by (14) and Cauchy-Schwarz, we have
@@
$\leq(8\left\|\sum_{\boldsymbol{\alpha}\in\mathbb{I}}\mathbf{B}_{\boldsymbol{\alpha}}\right\|+8n)\|\mathbf{x}\|_{2}^{2}+8\|\mathbf{x}\|_{1}^{2},$ (44)